INFRARED SPECTRA OF HYDRATED CLUSTERS OF GUANINE NUCLEOSIDES OBSERVED BY IR-UV DOUBLE RESONANCE SPECTROSCOPY

\underline{SHU-HEI URASHIMA}, HIROYA ASAMI, and HIROYUKI SAIGUSA, Graduate School of Arts and Sciences, Yokohama City University, Yokohama 236-0027, Japan.

Mono- and dihydrated clusters of guanosine and 2’-deoxyguanosine are produced by laser-desorption supersonic-jet cooling and their structures are identified by IR-UV double resonance spectroscopy combined with theoretical calculation. The results show that specific hydration structures around the sugar group exist in both nucleosides.^{a}^{b} It is also demonstrated that the dihydrated structures are strongly influenced by the presence or absence of the 2’-hydroxy group on the sugar.

aH. Saigusa, S. Urashima, and H. Asami, J. Phys. Chem. A, 2009, 113, 3455.

bH. Asami, S. Urashima, and H. Saigusa, Phys. Chem. Chem. Phys., 2009, 11, 10466.